# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision Tasks

### v2

@@ Abstract
We introduce intra-class memorability, where certain images within the same class are more memorable than others despite shared category characteristics. To investigate what features make one object instance more memorable than others, we design and conduct human behavior experiments, where participants are shown a series of images, and they must identify when the current image matches the image presented a few steps back in the sequence. To quantify memorability, we propose the Intra-Class Memorability score (ICMscore), a novel metric that incorporates the temporal intervals between repeated image presentations into its calculation. Furthermore, we curate the Intra-Class Memorability Dataset (ICMD), comprising over 2,500 images across five object classes with their ICMscores derived from 1,000 participants’ responses. Subsequently, we demonstrate the usefulness of ICMD by training AI models on this dataset for various downstream tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Surprisingly, high-ICMscore images impair AI performance in image recognition and continual learning tasks, while low-ICMscore images improve outcomes in these tasks. Additionally, we fine-tune a state-of-the-art image diffusion model on ICMD image pairs with and without masked semantic objects. The diffusion model can successfully manipulate image elements to enhance or reduce memorability. Our contributions open new pathways in understanding intra-class memorability by scrutinizing fine-grained visual features that result in the least and most memorable images and laying the groundwork for real-world applications in cognitive science and computer vision. All code, data, and models will be made publicly available.
Cognitive Science, Memorability
@@ 1 Introduction
Figure 1: An object instance can be more memorable than another within the same class. Intra-class memorability refers to the ability of instances within the same category to be remembered. First, we present example images with low and high memorability (first and second rows) for the exemplar flower class. A white pale daisy (Row 1) is less memorable than a striking violet osteospermum (Row 2). Images are identified as high or low memorability based on results from human behavior experiments. We then use these high-memorability and low-memorability images to train and test state-of-the-art AI models across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Each task is highlighted in dashed bubbles with distinct colors.
@@
As we experience the world, our memory selectively retains certain objects while overlooking others (McDermott & Roediger, 2018; Halamish & Stern, 2022; Leshikar et al., 2012). For instance, among numerous flowers we encounter, a pale white daisy might fade from memory, while a striking violet osteospermum with intricate shapes remains vividly etched (Fig. 1). What factors make some flowers more memorable than others? To answer this question, we first introduce intra-class memorability which refers to the ability of instances within the same object class to be remembered.
To delve into why certain items are more memorable than others, despite belonging to the same overall category, we conduct human behavior experiments using a continuous recognition task. In this task, participants are presented with a sequence of new and repeated images from the same class with delayed repetition and they are asked to identify those repeats. The proposed paradigm is novel compared to existing research, where continuous recognition tasks typically involve sequential presentations of images from different classes (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023; Hovhannisyan et al., 2021).
@@
To measure the memorability of individual instances, we introduce the Intra-Class Memorability score (ICMscore). Unlike prior evaluation metrics (Goetschalckx & Wagemans, 2019; Kramer et al., 2023; Isola et al., 2014; Akagunduz et al., 2019), our ICMscore incorporates average hit and miss rates for an item’s repetitions, weighted by the time intervals between its initial and repeated presentations. Missing a repeat after a short interval indicates low memorability (LM), while remembering a repeat after a long interval signifies high memorability (HM).
Due to the lack of intra-class memorability datasets, we contribute the Intra-Class Memorability dataset (ICMD). The dataset contains 5 classes spanning over 500 images per class with their corresponding ICMscore aggregated from the responses of 1000 participants in the human behavior experiments on continuous recognition tasks.
To demonstrate the effectiveness of our ICMD dataset, we train and test AI models on different subsets of ICMD images based on their ICMscores across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
@@
Our experimental results show that a memorability prediction model trained on ICMD captures representative features diagnostic for intra-class memorability. Moreover, surprisingly, while highly memorable images seem more salient and thus easier to remember, we found that these images actually impair AI performance in image recognition and continual learning tasks. Finally, we use images from ICMD with and without masked semantic objects to fine-tune the state-of-the-art image diffusion model (Brooks et al., 2022) using LoRA (Hu et al., 2021). The diffusion model can edit image elements to control memorability.
We highlight our key contributions:
@@
  • We introduce the novel and important problem of intra-class memorability and provide insights into why certain instances are more memorable than others, despite belonging to the same category.
@@
  • We challenge existing paradigms for collecting memorability data on naturalistic images by introducing a new human behavioral experiment on intra-class memorability. In addition, we propose a novel memorability metric, the ICMscore, which accounts for the time intervals before a repeated image is identified in continuous image recognition.
@@
  • We collect ICMscore from the responses of 1000 participants on 2500 images spanning across 5 object classes and contribute the first dataset on intra-class memorability, named as Intra-Class Memorability dataset (ICMD). To demonstrate the usefulness of our dataset, we introduce four computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
@@
  • Our results show that our memorability prediction model is capable of capturing key features for intra-class memorability. Surprisingly, while HM images are more salient and tend to be retained longer in memory, training AI models on these images can actually impair their performance in image recognition and continual learning tasks. Finally, trained on ICMD images with and without their semantic objects masked out, our fine-tuned state-of-the-art image diffusion model can enhance or reduce the memorability of a given image by manipulating their semantic regions.
@@ 2 Related Work
Research in cognitive psychology indicates that certain visual and contextual features—such as prominent visual characteristics, emotional valence, and semantic richness—substantially increase the likelihood of remembering specific objects over others (Schmidt, 1991). This implies that memorability is a predictable and universal property, rather than purely subjective (Konkle et al., 2010; Bylinskii et al., 2022; Isola et al., 2011). Based on this, Isola et al. (Isola et al., 2011) define “memorability” as the probability that a viewer will remember an image after a single viewing. Subsequent studies have investigated how intrinsic and extrinsic factors influence memorability. Extrinsic factors refer to an image’s categorical information, while intrinsic factors include other visual features beyond these categorical information (Khosla et al., 2015; Dubey et al., 2015; Hovhannisyan et al., 2021). However, none of these studies have effectively disentangled the effects of extrinsic and intrinsic features on memorability.
@@
To address this problem, subsequent works (Bylinskii et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023; Lu et al., 2020; Borkin et al., 2013) focus on specific super-categories, such as animals or indoor and outdoor scenes. In addition, some research has extended the studies on static images to videos, exploring memorability differences across video super-categories (Cohendet et al., 2018a; Newman et al., 2020; Cohendet et al., 2018b). Although these studies limit categories to reduce the impact of extrinsic factors on memorability, they still mix images from different sub-categories, making it difficult to conduct controlled studies that solely examine the effect of intrinsic features on memorability. To address this research gap, we introduce the concept of intra-class memorability. To investigate it, we propose a novel human behavior experimental paradigm to measure memorability differences among images within the same class.
@@
Several memorability datasets have been collected through human behavior experiments (Isola et al., 2011; Bylinskii et al., 2015; Dubey et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023), with LaMem (Khosla et al., 2015) as one of the largest for deep learning-based image memorability prediction. In these studies, participants typically view a sequence of images, identifying any repeated target images. Memorability scores are often calculated based on the proportion of participants who correctly identify these repeats (Isola et al., 2011; Bylinskii et al., 2015; Dubey et al., 2015). Some studies adjust scores to account for false alarms, where images are mistakenly identified as repeats (Bainbridge & Rissman, 2018; Goetschalckx & Wagemans, 2019; Kramer et al., 2023). The time interval between target presentations also influences memorability (Khosla et al., 2015), yet random sampling of intervals can lead to unfair comparisons due to inconsistent controls. In our experiment, we standardize time intervals and introduce a novel memorability score, named as Intra-Class Memorability score (ICMscore). It accounts for both hit rates and miss rates of repeated images across a predefined set of time intervals. Moreover, we contribute a new dataset, Intra-Class Memorability dataset (ICMD). It consists of 2500 images spanning five object classes with their corresponding ICMscores based on responses from over 1000 participants.
@@
Several studies have demonstrated that HM images can positively influence human cognitive states. For example, GANalyze (Goetschalckx et al., 2019) introduces a generative model that modifies image attributes to affect human emotional and cognitive responses. Peng et al. (Peng & Bainbridge, 2024) show that memorable images boost user engagement in image retrieval systems and improve recall accuracy in storytelling and educational contexts. Similarly, (Pataranutaporn et al., 2024) and (Siarohin et al., 2019) explore how AI-modified images and videos uniquely shape human memory. Thus, predicting memorability for a given image has become increasingly important (Lahrache & El Ouazzani, 2022; Squalli-Houssaini et al., 2018; Hagen & Espeseth, 2023). However, existing models for predicting memorability rely on both extrinsic and intrinsic features, making them unable to capture fine-grained memorability differences within the same class of images. In contrast, our experiments demonstrate that our model, trained on the ICMD dataset, outperforms existing predictors in accurately predicting intra-class memorability.
@@
Despite early efforts to explore the role of memorability in enhancing human cognition, its impact on computer vision tasks remains largely unexplored. Specifically, the connection between image features relevant to object recognition and those that explain memorability is still unclear. Understanding this relationship could have significant implications for fields like continual learning. Using our ICMD dataset, we present preliminary evidence that memorability affects AI performance in tasks such as image recognition and continual learning. To edit images with controls on memorability, we also build upon the existing generative models (Brooks et al., 2022; Hu et al., 2021) and introduce an image diffusion model, capable of manipulating image elements to enhance or reduce memorability.
@@ 3.1 Human Behavior Experiments
Different from the existing memorability datasets (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Hovhannisyan et al., 2021; Goetschalckx & Wagemans, 2019; Kramer et al., 2023), we focus on collecting a new Intra-Class Memorability dataset (ICMD), which features naturalistic images from five object classes and their associated human memorability responses. Fig. 2 presents the schematic of the human psychophysics experiments in the continuous object recognition tasks. In each trial, participants view a sequence of images, with both target and foil images belonging to the same category. Each image is followed by a blank screen to allow ample response time. Participants press the space bar when recognizing a previously seen image, with reaction times recorded during the 1,200 ms image presentation and the subsequent 1,600 ms blank screen. Before the start of an experiment, participants complete a guided demonstration to ensure task comprehension. During the experiment, immediate feedback follows each response: a green tick for correct answers and a red cross for incorrect ones. The image sequence is randomly shuffled, with target images reappearing after intervals of $t\in[8,16,24,32]$ images, uniformly distributed across these intervals. We determine these interval ranges through a pilot study exploring intervals $t$ from 2 to 64. The results revealed that intervals shorter than 8 led to near-ceiling performance, with participants achieving over 90% recall accuracy over all target images. Conversely, intervals exceeding 32 proved too challenging, with more than 30% of participants failing to identify any repeated images. Based on these findings, we selected the interval range of [8, 16, 24, 32] to ensure meaningful image-dependent variability in memory recall performance during the experiments. Each trial includes 100 unique target images, with foil images filling the sequence when targets are not shown, each target image and its repeat are shown only once to one participant. Sec. A.1 and Fig. A1 highlight differences in the experiment schematic between our study and previous works.
@@
We select 2,500 images from ImageNet (Deng et al., 2009) as target images for the experiment, evenly distributed across five categories: dogs, cars, flowers, jellyfish, and teapots. A total of 3,263 participants from diverse demographics are recruited on Amazon Mechanical Turk (MTurk) (Turk, 2012). To ensure data quality, participants are disqualified if they incorrectly identify more than three foil images as target images in a row or fail to respond within 60 seconds. After filtering the data from disqualified participants, we analyzed responses from the remaining 1,000 qualified participants, resulting in 2,000 trials and a total of 171,568 key presses, with an average of 86 key presses per trial. The average false positive rate of the key presses is 14.67%. All experiments are conducted with informed consent and follow protocols approved by our institution’s Institutional Review Board.
Figure 2: Schematic illustration of the human behavioral experiment on intra-class memorability. A sequence of new and repeated images is presented. Each image appears for 1200 ms, followed by a blank screen for 1600 ms. Participants are asked to identify those repeats by pressing the space on the keyboard. Red arrows indicate the correct identification of the repeated images. Unique foil images are inserted into the sequence whenever there is no target image in that location.
@@ 3.2 Intra-Class Memorability Score (ICMscore)
To assess intra-class memorability, we introduce the Intra-Class Memorability Score (ICMscore), which measures the memorability of an image independently of its categorical information. Previous research typically computes memorability scores using hit rates and miss rates of memory recall (Bainbridge & Rissman, 2018; Goetschalckx & Wagemans, 2019; Kramer et al., 2023). However, the interval $t$ between image repetitions significantly affects memorability (Khosla et al., 2015). As shown in Fig. A2 and Sec. A.2, memory accuracy declines with longer intervals due to memory decays and increasing cognitive loads over time (Wixted, 2004; Sikarwar & Zhang, 2024). Therefore, we incorporate the lengths of these time intervals into the ICMscore calculation. Next, we describe how we compute ICMscore for a target image $I$.
@@
For each target $I$, we receive key press responses from $n$ participants. The correctness of $i$th participant’s response is denoted as $x_{i}$, forming a response set $\{x_{1},x_{2},\dots,x_{n}\}$ for image $I$. A response of $x_{i}=1$ indicates a correct identification of a previously seen target, while $x_{i}=-1$ indicates an incorrect identification. The interval associated with each response $x_{i}$ is represented as $t_{i}$.
@@
Since human memory retention diminishes over longer intervals, a longer interval $t$ makes correct identification of seen images more challenging. Therefore, when a participant correctly identifies an image after a longer interval, we assign it a higher memorability score $s$ to reflect the increased difficulty associated with the larger $t$. For correct responses ($x_{i}=1$), the memorability score $s_{i}$ is computed using a linearly increasing function: $s_{i}=x_{i}\frac{t_{i}}{T}$, where $t_{i}$ is the interval between two presentations of a target image, and $T=32$ is the maximum interval. Conversely, as longer time intervals are more prone to incorrect identifications of repeated target images, errors during shorter intervals are more indicative of the image’s memorability. For incorrect responses ($x_{i}=-1$), shorter intervals incur a larger penalty, while longer intervals result in a smaller penalty. The memorability score for incorrect responses is given by $s_{i}=x_{i}\frac{T-t_{i}}{T}$.
The final ICMscore for each target image $I$ is calculated as the average of its memorability scores $s_{i}$ over all $n$ participants: ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$. We also analyzed ICMscore without the temporal penalty $t_{i}$ (Sec. A.6). Results indicate that without this penalty, ICMscore fails to capture intra-class memorability, eliminating its impact on downstream computer vision tasks.

### v3

@@ Abstract
We introduce intra-class memorability, where certain images within the same class are more memorable than others despite shared category characteristics. To investigate what features make one object instance more memorable than others, we design and conduct human behavior experiments, where participants are shown a series of images, and they must identify when the current image matches the image presented a few steps back in the sequence. To quantify memorability, we propose the Intra-Class Memorability score (ICMscore), a novel metric that incorporates the temporal intervals between repeated image presentations into its calculation. Furthermore, we curate the Intra-Class Memorability Dataset (ICMD), comprising over 5,000 images across ten object classes with their ICMscores derived from 2,000 participants’ responses. Subsequently, we demonstrate the usefulness of ICMD by training AI models on this dataset for various downstream tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Surprisingly, high-ICMscore images impair AI performance in image recognition and continual learning tasks, while low-ICMscore images improve outcomes in these tasks. Additionally, we fine-tune a state-of-the-art image diffusion model on ICMD image pairs with and without masked semantic objects. The diffusion model can successfully manipulate image elements to enhance or reduce memorability. Our contributions open new pathways in understanding intra-class memorability by scrutinizing fine-grained visual features behind the most and least memorable images and laying the groundwork for real-world applications in computer vision. We will release all code, data, and models publicly.
@@ 1 Introduction
Figure 1: An object instance can be more memorable than another within the same class. Intra-class memorability refers to the ability of instances within the same category to be remembered. First, we present example images for the five exemplar classes. Images are identified as low (Row 1) or high (Row 2) memorability based on results from human behavior experiments. We then use these high- and low-memorability images to train and test state-of-the-art AI models across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
@@
As we experience the world, our memory selectively retains certain objects while overlooking others [1, 2, 3]. For instance, among numerous flowers we encounter, a pale white daisy might fade from memory, while a striking violet osteospermum with intricate shapes remains vividly etched (Fig. 1). What factors make some flowers more memorable than others? To answer this question, we first introduce intra-class memorability which refers to the ability of instances within the same object class to be remembered. To delve into why certain items are more memorable than others, despite belonging to the same overall category, we conduct human behavior experiments using a continuous recognition task. In this task, participants are presented with a sequence of new and repeated images from the same class with delayed repetition and they are asked to identify those repeats. The proposed paradigm is novel compared to existing research, where continuous recognition tasks typically involve sequential presentations of images from different classes [4, 5, 6, 7, 8, 9].
@@
To measure the memorability of individual instances, we introduce the Intra-Class Memorability score (ICMscore). Unlike prior evaluation metrics [7, 8, 10, 11], our ICMscore incorporates average hit and miss rates for an item’s repetitions, weighted by the time intervals between its initial and repeated presentations. Missing a repeat after a short interval indicates low memorability (LM), while remembering a repeat after a long interval signifies high memorability (HM).
@@
Due to the lack of intra-class memorability datasets, we contribute the Intra-Class Memorability dataset (ICMD). The dataset contains 10 classes spanning over 500 images per class with their corresponding ICMscore aggregated from the responses of 2000 participants in the human behavior experiments on continuous recognition tasks. To demonstrate the effectiveness of our ICMD, we train and test AI models on different subsets of ICMD images based on their ICMscores across four downstream computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing. Our experimental results show that a memorability prediction model trained on ICMD captures representative features diagnostic for intra-class memorability. Moreover, surprisingly, while highly memorable images seem more salient and thus easier to remember, we found that these images actually impair AI performance in image recognition and continual learning tasks. Finally, we use images from ICMD with and without masked semantic objects to fine-tune the state-of-the-art image diffusion model [12] using LoRA [13] to control image memorability. We highlight our key contributions:
@@
1. We introduce the novel and important problem of intra-class memorability and provide insights into why certain instances are more memorable than others, despite belonging to the same category.
@@
2. We challenge existing paradigms for collecting memorability data on naturalistic images by introducing a new human behavior experiment on intra-class memorability. In addition, we propose a novel memorability metric, the ICMscore, which accounts for the time intervals before a repeated image is identified in continuous image recognition.
@@
3. We collect ICMscore from the responses of 2000 participants on 5000 images spanning across 10 object classes and contribute the first dataset on intra-class memorability, named as Intra-Class Memorability dataset (ICMD). To demonstrate the usefulness of our dataset, we introduce four computer vision tasks: memorability prediction, image recognition, continual learning, and memorability-controlled image editing.
@@
4. Our results show that our memorability prediction model is capable of capturing key features for intra-class memorability. Surprisingly, while HM images are more salient and tend to be retained longer in memory, training AI models on these images can actually impair their performance in image recognition and continual learning tasks. Finally, trained on ICMD images with and without their semantic objects masked out, our fine-tuned state-of-the-art image diffusion model can enhance or reduce the memorability of a given image by manipulating their semantic regions.
@@ 2 Related Work
Research in cognitive psychology indicates that certain visual and contextual features—such as prominent visual characteristics, emotional valence, and semantic richness—substantially increase the likelihood of remembering specific objects over others [14]. This implies that memorability is a predictable and universal property, rather than purely subjective [15, 16, 17]. Based on this, Isola et al. [17] define “memorability" as the probability that a viewer will remember an image after a single viewing. Subsequent studies have investigated how intrinsic and extrinsic factors influence memorability. Extrinsic factors refer to an image’s categorical information, while intrinsic factors include other visual features beyond these categorical information [5, 6, 9]. However, none of these studies have effectively disentangled the effects of extrinsic and intrinsic features on memorability.
@@
To address this problem, subsequent works [4, 7, 8, 18, 19] focus on specific super-categories, such as animals or indoor and outdoor scenes. In addition, some research has extended the studies on static images to videos, exploring memorability differences across video super-categories [20, 21, 22]. Although these studies limit categories to reduce the impact of extrinsic factors on memorability, they still mix images from different categories, making it difficult to conduct controlled studies that solely examine the effect of intrinsic features on memorability. To address this research gap, we introduce the concept of intra-class memorability. To investigate it, we propose a novel human behavior experimental paradigm focusing memorability among images within the same class.
@@
Several memorability datasets have been collected through human behavior experiments [17, 4, 6, 7, 8], with LaMem [5] as one of the largest for deep learning-based image memorability prediction. In these studies, participants typically view a sequence of images, identifying any repeated target images. Memorability scores are often calculated based on the proportion of participants who correctly identify these repeats [17, 4, 6]. Some studies adjust scores to account for false alarms, where images are mistakenly identified as repeats [23, 7, 8]. The time interval between target presentations also influences memorability [5], yet random sampling of intervals can lead to unfair comparisons due to inconsistent controls. In our experiment, we standardize time intervals and introduce a novel memorability score, named as Intra-Class Memorability score (ICMscore). It accounts for both hit rates and miss rates of repeated images across a predefined set of time intervals. Moreover, we contribute a new dataset, Intra-Class Memorability dataset (ICMD). It consists of 5000 images with their corresponding ICMscores based on responses from over 2000 participants.
@@
Several studies have demonstrated that HM images can positively influence human cognitive states. For example, GANalyze [24] introduces a generative model that modifies image attributes to affect human emotional and cognitive responses. Peng et al. [25] show that memorable images boost user engagement in image retrieval systems and improve recall accuracy in storytelling and educational contexts. Similarly, [26] and [27] explore how AI-modified images and videos uniquely shape human memory. Thus, predicting memorability for a given image has become increasingly important [28, 29, 30]. However, existing models for predicting memorability rely on both extrinsic and intrinsic features, making them unable to capture fine-grained memorability differences within the same class of images. In contrast, our experiments demonstrate that our model, trained on the ICMD, outperforms existing predictors in accurately predicting intra-class memorability.
@@
Despite early efforts to explore the role of memorability in enhancing human cognition, its impact on computer vision tasks remains largely unexplored. Specifically, the connection between image features relevant to object recognition and those that explain memorability is still unclear. Understanding this relationship could have significant implications for fields like continual learning. Using our ICMD, we present preliminary evidence that memorability affects AI performance in tasks such as image recognition and continual learning. To edit images with controls on memorability, we also build upon the existing generative models [12, 13] and introduce an image diffusion model, capable of manipulating image elements to enhance or reduce memorability.
@@ 3 Intra-Class Memorability Dataset(ICMD)
Figure 2: Schematic illustration of the human behavior experiment on intra-class memorability. A sequence of new and repeated images is presented. Each image appears for 1200 ms, followed by a blank screen for 1600 ms. Participants are asked to identify those repeats by pressing the space on the keyboard. Red arrows indicate the correct identification of the repeated images. Unique foil images are inserted into the sequence whenever there is no target image in that location.
@@ 3.1 Human Behavior Experiments
Different from the existing memorability datasets [4, 5, 6, 9, 7, 8], our work introduces a new Intra-Class Memorability Dataset (ICMD), comprising naturalistic images from ten distinct object categories along with corresponding human memorability responses. The selected categories—volcano, dog, phone, lemon, car, bird, palace, flower, jellyfish, and teapot—were chosen to ensure clear semantic separation with no category overlap. We curated a total of 5,000 target images from ImageNet [31], evenly distributed across these ten classes. Fig. 2 presents the schematic of the human psychophysics experiments in the continuous object recognition tasks. In each trial, participants view a sequence of images, with both target and foil images belonging to the same category. Each image is followed by a blank screen to allow ample response time. Participants press the space bar when recognizing a previously seen image, with reaction times recorded during the 1,200 ms image presentation and the subsequent 1,600 ms blank screen. Before the start of an experiment, participants complete a guided demonstration to ensure task comprehension. During the experiment, immediate feedback follows each response: a green tick for correct answers and a red cross for incorrect ones. The image sequence is randomly shuffled, with target images reappearing after intervals of $t\in[8,16,24,32]$ images, uniformly distributed across these intervals. We determine these interval ranges through a pilot study exploring intervals $t$ from 2 to 64. The results reveal that intervals shorter than 8 led to near-ceiling performance, with participants achieving over 93% recall accuracy over all target images. Conversely, intervals exceeding 32 proved too challenging, with more than 34% of participants failing to identify any repeated images. Based on these findings, we select the interval range of [8, 16, 24, 32] to ensure meaningful image-dependent variability in memory recall performance. Each trial includes 100 unique target images, with foil images filling the sequence when targets are not shown, each target image and its repeat are shown only once to one participant. Sec. A.1 and Fig. A1 highlight differences in the experiment schematic between our study and previous works.
@@
A total of 3,263 participants are recruited on Amazon Mechanical Turk (MTurk) [32]. To ensure data quality, participants are disqualified if they incorrectly identify more than three foil images as target images in a row or fail to respond within 60 seconds. After filtering the data from disqualified participants, we analyze responses from the remaining 2,000 qualified participants, resulting in 4,000 trials and a total of 346,537 key presses, with an average of 85 key presses per trial. The average false positive rate of the key presses is 15.24%. All experiments are conducted with informed consent and follow protocols approved by our institution’s Institutional Review Board.
@@ 3.2 Intra-Class Memorability Score (ICMscore)
To assess intra-class memorability, we introduce the Intra-Class Memorability Score (ICMscore), which measures the memorability of an image independently of its categorical information. Previous research typically computes memorability scores using hit rates and miss rates of memory recall [23, 7, 8]. However, the interval $t$ between image repetitions significantly affects memorability [5]. As shown in Fig. A2 and Sec. A.2, memory accuracy declines with longer intervals due to memory decays and increasing cognitive loads over time [33, 34]. Therefore, we incorporate the lengths of these time intervals into the ICMscore calculation.
@@
For each target $I$, we receive key press responses from $n$ participants. The correctness of $i$th participant’s response is denoted as $x_{i}$, forming a response set $\{x_{1},x_{2},\dots,x_{n}\}$ for image $I$. A response of $x_{i}=1$ indicates a correct identification, while $x_{i}=-1$ indicates an incorrect identification. The interval associated with each response $x_{i}$ is represented as $t_{i}$.
@@
Since human memory retention diminishes over longer intervals, a longer interval $t$ makes correct identification of seen images more challenging. Therefore, when a participant correctly identifies an image after a longer interval, we assign it a higher memorability score $s$ to reflect the increased difficulty associated with the larger $t$. For correct responses ($x_{i}=1$), the memorability score $s_{i}$ is computed using a linearly increasing function: $s_{i}=x_{i}\frac{t_{i}}{T}$, where $t_{i}$ is the interval between two presentations of a target image, and $T=32$ is the maximum interval. Conversely, as longer time intervals are more prone to incorrect identifications of repeated target images, errors during shorter intervals are more indicative of the image’s memorability. For incorrect responses ($x_{i}=-1$), shorter intervals incur a larger penalty, while longer intervals result in a smaller penalty. The memorability score for incorrect responses is given by $s_{i}=x_{i}\frac{T-t_{i}}{T}$. The final ICMscore for each target image $I$ is calculated as the average of its memorability scores $s_{i}$ over all $n$ participants: ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$.
To highlight the importance of the temporal penalty in our ICMscore formulation, we conducted an additional analysis by removing the temporal component $t_{i}$. See Sec. A.5 for details. Results show that without this penalty, ICMscore fails to capture intra-class memorability effectively, diminishing its utility for downstream computer vision tasks. Furthermore, to demonstrate that ICMscore is not merely a variant of existing memorability metrics defined in [23, 7, 8], we applied an existing memorability prediction model trained on LaMem dataset [5] to estimate memorability scores for images in our ICMD dataset. The two sets of scores show only a weak positive correlation, with a Pearson coefficient of 0.26. This suggests that ICMscore captures distinct memorability-related features overlooked by prior metrics and serves a complementary role in memorability evaluation.
Figure 3: Correlation analysis and Grad-CAM visualization. (a) Pearson correlation between feature distances of images to their categorical centroids and their ICMscore. The y-axis shows correlation values while the x-axis displays class labels. Correlations range from $r=0.21$ for “Flower" to $r=0.21$ for “Phone" ($p<0.05$ for all ten classes). (b) Grad-CAM visualizations displaying importance maps for three high memorability (HM) images. These visualizations highlight key regions by computing gradients of the memorability predictor with respect to feature maps to identify areas contributing most to memorability scores. For HM images, features enhancing memorability are highlighted according to the color scale shown in the right color bar. The first column presents original images, the second column shows visualizations from the model trained on LaMem dataset [5], and the third column displays results from our model trained on ICMD.